### v1

@@ 1 Introduction
This paper summarizes the LNQ 2023 challenge and is structured as follows. First, a review of existing datasets used to perform lymph node quantification is presented in Section 2. Then, the design of the LNQ challenge is given in Section 3. Section 4 presents the evaluation strategy of the challenge (metrics and ranking scheme). Participating methods are then described and compared in Section 5. Finally, Section 6 presents the results of the participating team, and Section 7 provides a discussion and concludes the paper.
@@ 2 Related Works
On the methodological side, a wide range of weakly-supervised methods have been proposed. Nonetheless, a thorough review of weakly-supervised methodologies is out of the scope of this paper, except to note that it it an active area of research with many unsolved challenges. We refer the interested reader to Zhang et al. (2020) for a recent review of these.
@@
Hodgkin’s lymphoma ($7\%$) Hodgkin’s lymphoma ($5\%$) Hodgkin’s lymphoma ($8\%$)
@@ 3.2.1 Data overview
The challenge cohort is a cross-institutional dataset of 513 chest CT scans acquired during patient treatment for various cancer types. The dataset originates from the Tumor Imaging Metrics Core (TIMC), a multi-institutional imaging core lab (Massachusetts General Hospital, Dana Farber Cancer Institute, Brigham and Women’s Hospital, Boston, MA, USA) that provides multimodality imaging measurements to evaluate treatment response in patients enrolled in oncology clinical trials. This dataset was used for training, validation, and testing. Figure 1 presents an overview of the challenge dataset.
@@
Table 1 shows the distribution of features (patient sex, cancer, slice thickness, and in-plane resolution) between the training, validation, and test sets.
@@ 3.2.3 Annotation protocol
Volumetric segmentations were performed using 3D Slicer (Pieper et al., 2004; Fedorov et al., 2012). All lymph nodes were segmented by a trained radiologist (TI), with the support of an instructor in radiology (BS), and in consultation with a senior radiologist (HJ). For each patient, the CT scan and the bidimensional measurements of the target lymph node were loaded into 3D Slicer (version 4.11). The Segment Editor module was then used to manually delineate the lymph node boundary at the native resolution of the CT scan. The Draw tool within the Editor module was employed to draw freehand boundaries on axial cross-sections while referencing the sagittal and coronal planes.
@@ 3.2.4 Data curation
Images and segmentation masks were distributed as NRRD files (.nrrd). The training and validation data was available via the challenge page  [^1] https://lnq2023.grand-challenge.org/data/. As we expect this dataset to be used for other purposes, the data has been accepted and will be be released soon on the Cancer Imaging Archive (TCIA).
@@ 5 Participating methods
In this section, we summarize the methods used by these 6 teams. Each method is assigned a unique color code used in the tables and figures.
@@
This team employed a semi- and weakly-supervised approach. The authors proposed incorporating the TotalSegmentator toolbox (Wasserthal et al., 2023) to generate pseudo labels and loss masking for handling incomplete annotations in combination with supervised learning using the nnUNet framework (Isensee et al., 2021). Alongside the challenge training data, the team utilized the fully annotated TCIA CT Lymph Nodes dataset (Roth et al., 2015), providing annotations of all pathologic lymph nodes. Instead of training solely on pathologic lymph nodes, they replaced original annotations with refined annotations containing all visible lymph nodes (Bouget et al., 2023). Furthermore, to increase the dataset size, the authors included the public lung cancer datasets NSCLC radiomics (Aerts et al., 2014), NSCLC radiogenomics (Bakr et al., 2018), and NSCLC interobserver (Wee et al., 2019). First, they created ROIs by cropping each volume to the lung bounding box via the TotalSegmentator. To handle incomplete annotations, pseudo labels were generated from TotalSegmentator structures. Those structures should, by definition, not contain any lymph nodes. Labeling those structures as background increased the training supervision significantly. The authors also masked the loss of remaining unlabeled voxels from the training process. Only one nnUNet instance was trained on the preprocessed data with adjusted hyperparameters for learning rate and intensity clipping. Furthermore, the authors applied a postprocessing step on the segmentation output by discarding all lymph node components with a shortest axis diameter less than 10 mm. In their final challenge report (Fischer et al., 2024), the authors show that this postprocessing hurt the overall performance and was based on their misinterpreting the challenge goal.
@@
The team proposed an ensemble of five segmentation models based on full-resolution nnUNet trained on multiple anatomical priors as additional input (Isensee et al., 2021; Engelson et al., 2024b). To address the challenges arising from the weak annotations of the training data, a probabilistic lymph node atlas was registered to the training data to identify regions with high lymph node occurrence, which was then used for loss weighting and post-processing. Further, the authors addressed the heterogeneity of the training data and lymph node appearance by using a strong augmentation called GIN &\&& IPA introduced by Ouyang et al. (2021). Preprocessing contained the generation of the anatomical priors using atlas-to-patient registration based on selected segmentation masks from the TotalSegmentator algorithm (Wasserthal et al., 2023), cropping to the lung region, and normalization. For post-processing, the threshold for binarization was lowered according to the lymph node atlas, and segmentation masks below a minimum diameter size of 5 mm were removed. The input data consisted of the weakly-labeled LNQ 2023 training data and publicly available CT Lymph Node dataset (Roth et al., 2015) with refined annotations Bouget et al. (2023) and the St. Olavs Hospital dataset (Bouget et al., 2019). The fully annotated data was oversampled during training. After the publication of the fully annotated LNQ 2023 validation and test data, the authors realized that the restriction to a minimum lymph node size degraded segmentation accuracy and that refining on the fully annotated datasets improved performance, as mentioned in their report (Engelson et al., 2024a).
Table 2: Metrics values and corresponding scores of submission. Median and interquartile values are presented. The best results are given in bold. Arrows indicate the favourable direction of each metric.
@@
The team used encoder-decoder type deep neural networks. Although the input images are 3D volumes, they use a 2D deep neural network model with multi-slice inputs (2.5D CNNs). They used U-Net as the basic segmentation network and replaced its encoder part with EfficientNet (Tan and Le, 2019). In their model, N consecutive slices in each input volume were concatenated and treated as an N channel 2D input image. The input image was processed to produce one slice in a segmentation mask volume. The input slices were along the transverse plane. In the training phase, slices were randomly selected M times from each volume in the training dataset in each epoch. They used a sum of Dice loss and cross-entropy loss as the loss function. The weights between the two losses were the same. They divided the provided dataset into training and validation data. The numbers of training and validation samples were 353 and 37 in the dataset, respectively. The optimizer was AdamW and the learning rate was decreased at every epoch with cosine annealing. The number of epochs was 100, and they used the model with the lowest loss value for the validation data as the final model. Random intensity shift and random affine transformation were performed as the augmentation. They conducted hyper-parameter tuning. The hyper-parameters included the encoder size, the number of slices, and the initial learning rate. As the results of hyper-parameter tuning, they selected EfficientNet-B7 as the encoder and 5 as the number slices. The initial learning rates were chosen as $10^{-3}$. In the inference phase, a volume in each case was processed slice by slice. They selected the largest connected component as post-processing step.
@@ 6.2 Analysis of the variations in performance
To determine if certain patient attributes are predictive of segmentation difficulty, we analyzed the distribution of the best Dice scores across all methods for each case, considering key patient attributes such as cancer type and sex. The results are shown in Figure 4. We found that the best Dice scores had a similar distribution across sexes, as depicted in Figure 4a. Conversely, Figure 4b shows the distribution of scores by patient cancer type. It can be observed that patients with Hodgkin’s lymphoma had statistically lower scores compared to those with chronic lymphocytic leukemia (CLL) ($p=0.009$), renal cell carcinoma ($p=0.003$), breast cancer ($p=0.008$), and other types ($p=0.27$). This could be attributed to the fact that lymph nodes associated with Hodgkin’s lymphoma are typically bulky and organized in conglomerates, making their segmentation more challenging.
@@ 7.1 Full supervision leads to higher performance
The top-ranked teams all utilized existing fully annotated datasets to train their frameworks. Notably, the top two teams first trained their models without the weakly annotated LNQ data, subsequently refining their models using partial annotations. In contrast, methods that solely leveraged weak annotations underperformed compared to these fully supervised approaches. This underscores that full supervision still outperforms existing weakly supervised approaches, even when using a smaller amount of training data.
@@ 7.2 Limitations
First, there was a noticeable shift in the distribution of cancer types between the training, validation, and testing sets. Despite this gap, similar performance levels were observed across all cancer types, except for Hodgkin’s lymphoma. Importantly, the proportion of Hodgkin’s lymphoma cases was consistent across all three sets, suggesting that the lower performance for this type was due to the inherent difficulty in segmenting its bulky and conglomerated lymph nodes rather than the distribution gap.
@@ 7.3 Conclusion
This work was supported by the National Institutes of Health (R01EB032387, P41EB015902, P41EB028741, R01CA235589, 5P30CA006516, 1U24CA258511), National Cancer Data Ecosystem (Task Order No. 413 HHSN26110071 under Contract No. HHSN261201500003l) and EOSS5 and EOSS3 Diversity and Inclusion grants from Chan-Zuckerberg Initiative.

### v2

@@ 1 Introduction
This paper summarizes the LNQ 2023 challenge and is structured as follows. First, a review of existing datasets used to perform lymph node quantification is presented in Section 2. Then, the design of the LNQ challenge is given in Section 3. Section 4 presents the evaluation strategy of the challenge (metrics and ranking scheme). Participating methods are then described and compared in Section 5. Finally, Section 6 presents the results of the participating teams, and Section 7 provides a discussion and concludes the paper.
@@ 2 Related Works
On the methodological side, a wide range of weakly-supervised methods have been proposed. Nonetheless, a thorough review of weakly-supervised methodologies is out of the scope of this paper, except to note that it it an active area of research with many unsolved challenges. We refer the interested reader to Zhang et al. (2020) for a review of these.
@@
Hodgkin lymphoma ($7\%$) Hodgkin lymphoma ($5\%$) Hodgkin lymphoma ($8\%$)
@@ 3.2.1 Data overview
The challenge cohort is a cross-institutional dataset of 513 chest CT scans acquired during patient treatment for various cancer types. The dataset originates from the Tumor Imaging Metrics Core (TIMC), a multi-institutional imaging core lab (Massachusetts General Hospital, Dana-Farber Cancer Institute, Brigham and Women’s Hospital, Boston, MA, USA) that provides multimodality imaging measurements to evaluate treatment response in patients enrolled in oncology clinical trials. This dataset was used for training, validation, and testing. Figure 1 presents an overview of the challenge dataset.
@@
The dataset was randomly split into training ($76\%$), validation ($4\%$) and test ($20\%$) sets. Table 1 shows the distribution of features (patient sex, cancer, slice thickness, and in-plane resolution) across these sets.
@@ 3.2.3 Annotation protocol
Volumetric segmentations were performed using 3D Slicer (Pieper et al., 2004; Fedorov et al., 2012). All lymph nodes were segmented by a trained radiologist (TI), with the support of an instructor in radiology (BS), and in consultation with a senior imaging specialist (HJ). For each patient, the CT scan and the bidimensional measurements of the target lymph node were loaded into 3D Slicer (version 4.11). The Segment Editor module was then used to manually delineate the lymph node boundary at the native resolution of the CT scan. The Draw tool within the Editor module was employed to draw freehand boundaries on axial cross-sections while referencing the sagittal and coronal planes.
@@ 3.2.4 Data curation
Images and segmentation masks were distributed as NRRD files (.nrrd). The training and validation datasets were made publicly available via the challenge page  [^1] https://lnq2023.grand-challenge.org/data/. In contrast, the testing set test set remained private. As we expect this annotated test dataset to be used for other purposes, the data has been on the Cancer Imaging Archive (TCIA) (Idris et al., 2024).
@@ 4.2 Ranking scheme
Table 2: Comparisons of the proposed techniques in terms of methodology and implementation details.
Skeleton Suns IMR CompAI Hilab IMI SKJP
Network architecture 3D U-Net (nnU-Net) 3D U-Net (nnU-Net) 3D U-Net (nnU-Net) 3D U-Net (nnU-Net) 3D V-Net (nnU-Net) 2.5D U-Net
TCIA CT Lymph Nodes + TCIA CT Lymph Nodes + TCIA CT Lymph Nodes + TCIA CT Lymph Nodes +
Use of external Bouget refinements Bouget refinements Bouget refinements Bouget refinements
datasets St. Olavs Hospital St. Olavs Hospital
NSCLC datasets $\times$ $\times$
Cropping $\times$ Lung and airway Lung Lung Lung $\times$
Weakly Supervision Background Background Partial-supervised Probabilistic
Approach pseudo-labeling Self-supervision pseudo-labeling loss functions atlas $\times$
$\times$ Small component Small component Small component Small component Largest component
Post-processing removal removal removal removal selection
@@ 5 Participating methods
In this section, we summarize the methods used by these 6 teams. Each method is assigned a unique color code used in the tables and figures. Brief comparisons of the proposed techniques in terms of methodology and implementation details are presented in Table 2.
@@
This team employed a semi- and weakly-supervised approach. The authors proposed incorporating the TotalSegmentator toolbox (Wasserthal et al., 2023) to generate pseudo labels and loss masking for handling incomplete annotations in combination with supervised learning using the nnU-Net framework (Isensee et al., 2021). Alongside the challenge training data, the team utilized the fully annotated TCIA CT Lymph Nodes dataset (Roth et al., 2015), providing annotations of all pathologic lymph nodes. Instead of training solely on pathologic lymph nodes, they replaced original annotations with refined annotations containing all visible lymph nodes (Bouget et al., 2023). Furthermore, to increase the dataset size, the authors included the public lung cancer datasets NSCLC radiomics (Aerts et al., 2014), NSCLC radiogenomics (Bakr et al., 2018), and NSCLC interobserver (Wee et al., 2019). First, they created ROIs by cropping each volume to the lung bounding box via the TotalSegmentator. To handle incomplete annotations, pseudo labels were generated from TotalSegmentator structures. Those structures should, by definition, not contain any lymph nodes. Labeling those structures as background increased the training supervision significantly. The authors also masked the loss of remaining unlabeled voxels from the training process. Only one nnU-Net instance was trained on the preprocessed data with adjusted hyperparameters for learning rate and intensity clipping. Furthermore, the authors applied a postprocessing step on the segmentation output by discarding all lymph node components with a shortest axis diameter less than 10 mm. In their final challenge report (Fischer et al., 2024), the authors show that this postprocessing hurt the overall performance and was based on their misinterpreting the challenge goal.
@@
The team proposed an ensemble of five segmentation models based on full-resolution nnU-Net trained on multiple anatomical priors as additional input (Isensee et al., 2021; Engelson et al., 2024b). To address the challenges arising from the weak annotations of the training data, a probabilistic lymph node atlas was registered to the training data to identify regions with high lymph node occurrence, which was then used for loss weighting and post-processing. Further, the authors addressed the heterogeneity of the training data and lymph node appearance by using a strong augmentation called GIN &\&& IPA introduced by Ouyang et al. (2021). Preprocessing contained the generation of the anatomical priors using atlas-to-patient registration based on selected segmentation masks from the TotalSegmentator algorithm (Wasserthal et al., 2023), cropping to the lung region, and normalization. For post-processing, the threshold for binarization was lowered according to the lymph node atlas, and segmentation masks below a minimum diameter size of 5 mm were removed. The input data consisted of the weakly-labeled LNQ 2023 training data and publicly available CT Lymph Node dataset (Roth et al., 2015) with refined annotations Bouget et al. (2023) and the St. Olavs Hospital dataset (Bouget et al., 2019). The fully annotated data was oversampled during training. After the publication of the fully annotated LNQ 2023 validation and test data, the authors realized that the restriction to a minimum lymph node size degraded segmentation accuracy and that refining on the fully annotated datasets improved performance, as mentioned in their report (Engelson et al., 2024a).
Table 3: Metrics values and corresponding scores of submission. Median and interquartile values are presented. The best results are given in bold. Arrows indicate the favorable direction of each metric.
@@
The team used encoder-decoder type deep neural networks. Although the input images are 3D volumes, they use a 2D deep neural network model with multi-slice inputs (2.5D CNNs). Specifically, they used a 2D U-Net as the segmentation network and replaced its encoder part with EfficientNet (Tan and Le, 2019). In their model, N consecutive slices in each input volume were concatenated and treated as an N channel 2D input image. The input image was then processed to produce one slice in a segmentation mask volume. The input slices were along the transverse plane. During training, slices were randomly selected M times from each volume in the training dataset in each epoch. The loss function was a sum of Dice loss and cross-entropy loss with equal weights. The dataset was split into 353 training samples and 37 validation samples. The optimizer was AdamW and the learning rate was decreased at every epoch with cosine annealing. The model was trained for 100 epochs, and the version with the lowest validation loss was selected as the final model. Random intensity shifts and random affine transformations were applied as data augmentation. Hyperparameter tuning was conducted, including adjustments to the encoder size, number of slices, and initial learning rate. As a result, EfficientNet-B7 was chosen as the encoder, with 5 slices and an initial learning rate of $10^{-3}$. During inference, each volume was processed slice by slice, and post-processing involved selecting the largest connected component.
@@ 6.2 Analysis of the variations in performance
To determine if certain patient attributes are predictive of segmentation difficulty, we analyzed the distribution of the best Dice scores across all methods for each case, considering key patient attributes such as cancer type and sex. The results are shown in Figure 4. We found that the best Dice scores had a similar distribution across sexes, as depicted in Figure 4a. Conversely, Figure 4b shows the distribution of scores by patient cancer type. It can be observed that patients with Hodgkin lymphoma had statistically lower scores compared to those with chronic lymphocytic leukemia (CLL) ($p=0.009$), renal cell carcinoma ($p=0.003$), breast cancer ($p=0.008$), and other types ($p=0.27$). This could be attributed to the fact that lymph nodes associated with Hodgkin lymphoma are typically bulky and organized in conglomerates, making their segmentation more challenging.
@@ 7.1 Full supervision leads to higher performance
The top-ranked teams all utilized existing fully annotated datasets to train their frameworks. Notably, the top two teams first trained their models without the weakly annotated LNQ data, subsequently refining their models using partial annotations. TotalSegmentator was especially used to remove false positives. In contrast, methods that solely leveraged weak annotations underperformed compared to these fully supervised approaches. This underscores that full supervision still outperforms existing weakly supervised approaches, even when using a smaller amount of training data.
@@ 7.2 Limitations
First, there was a noticeable shift in the distribution of cancer types between the training, validation, and testing sets. Despite this gap, similar performance levels were observed across all cancer types, except for Hodgkin lymphoma. Importantly, the proportion of Hodgkin lymphoma cases was consistent across all three sets, suggesting that the lower performance for this type was due to the inherent difficulty in segmenting its bulky and conglomerated lymph nodes rather than the distribution gap.
@@ 7.3 Conclusion
This work was supported by the National Institutes of Health (R01EB032387, P41EB015902, P41EB028741, R01CA235589, 5P30CA006516, 1U24CA258511), National Cancer Data Ecosystem (Task Order No. 413 HHSN26110071 under Contract No. HHSN261201500003l) and EOSS5 and EOSS3 Diversity and Inclusion grants from Chan-Zuckerberg Initiative. R.D. received a Marie Skłodowska-Curie fellowship No 101154248 (project: SafeREG).